The stock price of Google is $587. You have $10,000 to invest. The monthly interest rate is 0.5%.

a. You think the stock price will go down soon, and want to trade 20 shares. What should you do? Enter 20 for buying 20 shares (on margin if necessary), or -20 for selling or short-selling 20 shares.

b. If the initial margin is 50%, what is the minimum additional dollar amount that you have to deposit in your brokerage account?

c. What is your initial percentage margin (entered as a decimal number)?

d. Two months later, the stock price is $607. Google paid a dividend of $8 per share just before the two months were over. What is your percentage margin (entered as a decimal number)?
Answer #1

a) Since the view is that stock price would decrease, so corrct answer is -20.

b) Value = 50%*587*20 = 5870

c) Inital margin = 0.50

d) Two months later gain = 20*20 + 20*8 = 560

So margin = (5870 + 560 )/ (587*20 + 560) = 52.28%